OpenStudy (anonymous):

Find all unit vectors perpendicular to the surface x^2 + 2yz^3 = 3 at the point (1, -1, -1)

OpenStudy (anonymous):

find gradient and then neg. gradient should be all i believe, if I am not mistaken

OpenStudy (anonymous):

then once you find those, divide by length

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[f(x,y,z)=x^2+2yz^3-3=0 \rightarrow \nabla f=(2x,2z^3,0)\]will be normal to the surface at (x,y,z). Then,\[\pm \frac{\nabla f(x,y,z)}{|\nabla f(x,y,z)|}=\pm \frac{(2x,2z^3,0)}{\sqrt{4x^2+4z^6}}\]

OpenStudy (anonymous):

Substitute \[(x,y,z)=(1,-1,-1)\]

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

forgot z

OpenStudy (anonymous):

why the 0 term at the end?

OpenStudy (anonymous):

yeah...because i just got up ;)

OpenStudy (anonymous):

oh, ok I see

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

yeah, that is the gradient I got, with the last term being 6yz^2

OpenStudy (anonymous):

\[\nabla (x,y,z)=(2x,2z^3,6yz^2)\]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

how do you type in all those commands? that is cool, is there a manual around that I could look at some of those in?

OpenStudy (anonymous):

\[\pm \frac{\nabla (x,y,z)}{|\nabla (x,y,z)|}= \pm \frac{(2x,2z^3,6yz^2)}{\sqrt{4x^2+4z^6+36y^2z^4}}\]

OpenStudy (anonymous):

I wish there was a manual. I find everything by trial-and-error. Just use the equation editor on the bottom.

OpenStudy (anonymous):

You can take out a common factor of 2 between the numerator and denominator if you're desperate...

OpenStudy (anonymous):

But, cquinn, all you need to do now is sub (x,y,z)=(1,-1,-1).

OpenStudy (anonymous):

ahh, didn't recognize that at the bottom. will fiddle around with it some

OpenStudy (anonymous):

have fun!

OpenStudy (anonymous):

Thanks so much!! This helps a TON

OpenStudy (anonymous):

np

OpenStudy (anonymous):

fan us ;)